# Uncovering Latent Arguments in Social Media Messaging by Employing LLMs-in-the-Loop Strategy

### v1

@@ A.2 Prompt Templates
A.3 Resulting Arguments
@@ A.3.1 Climate Campaigns
Economy_pro ”Against natural gas bans due to their potential economic harm to families and workers.”, ”Joining Energy Citizens to support the natural gas and oil industry’s economic impact and sustainability efforts.”, ”Acknowledging and supporting the oil and natural gas industry’s critical role in economic growth and job creation.”, ”Permian Basin’s significant contribution to the energy sector’s growth and the positive economic impact of increased drilling activity.”, ”Considering the economic and community impact of taxation on the energy sector, emphasizing support for energy-related jobs and services.”, ”The Line 3 Replacement project’s role in boosting Minnesota’s economy through job creation, supporting Tribal communities, and enhancing safety, while positively impacting the environment.”, ”Recognizing the oil and gas industry’s vital role in economic growth, while addressing challenges like automation and training the future workforce.”
@@
ClimateSolution ”Adopting propane as a versatile and environmentally friendly energy solution to support climate change mitigation efforts.”, ”Leveraging innovative technologies and the synergy between natural gas and renewables as key strategies for climate change mitigation.”, ”Leveraging diverse, innovative, and market-based strategies to effectively address climate change and promote environmental sustainability.”, ”Adoption of RGGI in Pennsylvania, international cooperation on climate action, and exploring market-based, conservative solutions for sustainable economic and environmental health.”, ”Investing in innovative technologies like ICHOR’s RKI Injector and AT&T’s initiatives to drastically reduce emissions and improve fuel efficiency.”
@@
Pragmatism ”Pragmatic, multi-tiered action on climate change, emphasizing renewable energy investments and policy for sustainable development and job growth.”, ”Practical investments in sustainable infrastructure and energy diversity to address climate change and ensure future resilience.”, ”Pragmatic approach to energy transition, emphasizing the need to harmonize environmental sustainability, economic growth, and job preservation.”, ”Recognizing and utilizing natural gas as a pragmatic, efficient energy solution for reliability and affordability.”, ”Pragmatic approach to energy policy that balances clean energy transition with economic growth and job preservation.”, ”Pragmatic approach to energy, balancing economic growth and environmental protection through Colorado’s oil and gas industry.”, ”Pragmatic approach to energy, emphasizing the need for oil and safe pipeline infrastructure for global energy security.”, ”Pragmatic policy changes and discussions that address workers’ rights, environmental sustainability, and societal challenges for improvement and growth.”, ”Pragmatic solutions to environmental challenges through technological innovation, collaboration, and strategic policy efforts to ensure sustainability and security.”
@@
Patriotism ”Prioritizing domestic energy production as a patriotic duty to ensure America’s independence, security, and economic prosperity.”, ”Maintaining local oil and gas production to ensure economic and energy security, highlighting patriotism through self-reliance.”, ”Supporting the oil and gas industry as a patriotic duty, emphasizing its critical role in national security and economic health.”, ”Supporting natural gas to ensure energy independence and consumer choice, highlighting patriotism through energy security and freedom.”, ”Rallying around patriotism, energy independence, and conservative values to protect and promote American interests and security.”, ”Maintaining local oil and gas production to ensure California’s energy independence and protect consumers from rising fuel costs.”, ”Energy independence for Israel through faith-inspired and scientifically supported oil and gas exploration, and diverse leadership in community engagement.”
@@
AgainstClimatePolicy ”Renewable energy investment as a strategy for economic growth and job creation while addressing climate change.”, ”Transitioning to a clean energy economy, emphasizing job creation, sustainability, and combating climate change for economic recovery.”, ”Significant investment in clean energy infrastructure to create jobs and power homes sustainably, driving economic and environmental progress.”, ”Passage of the bipartisan infrastructure deal to enhance clean water, transit, and create numerous good jobs.”, ”American Jobs Plan as a catalyst for economic revitalization through clean energy, creating jobs, and improving infrastructure.”, ”Enhancing sustainability through clean energy initiatives, education, infrastructure improvement, and worker health and safety in the clean energy sector.”, ”Local and state-level initiatives that promote clean energy, sustainability, and environmental protection to build a more sustainable future.”, ”Comprehensive community and environmental betterment through sustainable economic practices, fair wages, and innovative climate solutions.”, ”Rigorous scrutiny and equitable policies in clean-energy projects to ensure they genuinely benefit the environment and communities.”
@@
Economy_clean ”Renewable energy investment as a strategy for economic growth and job creation while addressing climate change.”, ”Transitioning to a clean energy economy, emphasizing job creation, sustainability, and combating climate change for economic recovery.”, ”Significant investment in clean energy infrastructure to create jobs and power homes sustainably, driving economic and environmental progress.”, ”Passage of the bipartisan infrastructure deal to enhance clean water, transit, and create numerous good jobs.”, ”American Jobs Plan as a catalyst for economic revitalization through clean energy, creating jobs, and improving infrastructure.”, ”Enhancing sustainability through clean energy initiatives, education, infrastructure improvement, and worker health and safety in the clean energy sector.”, ”Local and state-level initiatives that promote clean energy, sustainability, and environmental protection to build a more sustainable future.”, ”Comprehensive community and environmental betterment through sustainable economic practices, fair wages, and innovative climate solutions.”, ”Rigorous scrutiny and equitable policies in clean-energy projects to ensure they genuinely benefit the environment and communities.”
@@
HumanHealth ”Reducing air pollution from oil and gas industries to protect children’s health and ensure environmental justice.”, ”Electing knowledgeable leaders and supporting organizations to address climate change’s significant health and humanitarian impacts.”, ”Climate change’s health impacts through informed initiatives and equity-focused efforts to protect vulnerable communities.”, ”Leadership that understands and addresses the health impacts of climate change, emphasizing immediate action for global health and stability.”, ”Environmental health risks, from fracking and gas stoves to substance abuse, to protect and improve public health.”, ”Combining environmental activism with health equity to combat pollution and support sustainable, just communities.”, ”Awareness of gas stoves’ health risks, promoting safer alternatives, and implementing preventive measures to protect indoor air quality and prevent asthma.”, ”Informed, compassionate leadership and global medical support to address healthcare challenges, ensuring dignity and comprehensive care across life stages and crises.”, ”Addressing and mitigating air pollution sources to protect public health and promote environmental sustainability.”, ”Proactive engagement and informed responses to health and safety challenges exacerbated by environmental changes and societal issues.”
@@
FutureGeneration ”Urgent environmental action and policy change to protect future generations from the health impacts of climate change.”, ”Educating children about climate change through open discussions to inspire future generations to take action for their planet’s health.”, ”Collective action led by motivated parents to combat climate change for a healthier future for children.”, ”Empowering the youth to lead climate change solutions, emphasizing education, activism, and support for vulnerable communities.”, ”Civic engagement and local policy reforms to secure a safer, fairer, and sustainable future for the next generation.”, ”Sustainable and renewable energy initiatives led by Indigenous and community efforts to ensure a healthier planet for future generations.”, ”Integrated solutions to climate change that also address social, racial, and economic inequities for future generations’ benefit.”, ”Sustainable energy and community initiatives that create educational, economic, and environmental benefits for current and future generations.”
@@
Environmental ”Ending fossil fuel use, implementing clean energy transitions, and enacting protective regulations to combat climate change and pollution.”, ”Urgent legislative action on climate change to enhance resiliency against extreme weather events and secure a sustainable future.”, ”Collective action and education to combat climate change, emphasizing community involvement and urgent policy measures.”, ”Community resilience through local initiatives, transitioning to clean energy, sustainable practices, and comprehensive support for businesses impacted by environmental disasters.”, ”Environmental protection and Indigenous rights by opposing oil pipelines that threaten natural resources and exacerbate the MMIW epidemic.”, ”Vigilance and community action against organized crime targeting green energy resources, and promotes sustainable living practices.”, ”Proactive climate change solutions through innovative forestry, ecosystem restoration projects, clean energy investment, and urgent policy actions to safeguard the future.”, ”Climate change’s role in exacerbating droughts, fires, and heatwaves, underscored by scientific consensus and observable impacts like Minnesota’s rising temperatures.”, ”Active participation in environmental conservation efforts led by Derek Kilmer to restore Puget Sound, protect wilderness, and combat climate change.”
@@
Animals ”Re-protection of gray wolves, recognizing their vital ecological role, and calls for increased conservation efforts and public support.”, ”Urgent financial contributions, leveraging donation matches to combat threats to endangered wildlife like habitat loss and climate change.”, ”Urgent climate action to protect Alaska’s fisheries, crucial for both the economy and marine ecosystems.”, ”Continued investment in climate and clean energy solutions to protect birds from climate change impacts.”, ”Immediate public action to conserve wildlife and address climate change’s role in accelerating species extinction.”, ”Habitat restoration and conservation to support wildlife, enhance outdoor activities, and address climate change for future ecological balance.”, ”Protecting and restoring ocean ecosystems to preserve marine biodiversity and combat climate change’s impacts on animals.”, ”Awareness and action against the environmental and ethical harms of animal agriculture, emphasizing sustainable choices and animal rights.”, ”Habitat conservation, creating pollinator gardens, and innovative habitat restoration to protect bats, bees, and wildlife crucial for ecosystem health and biodiversity.”, ”Promoting renewable energy projects as a means to generate clean power, support local economies, and protect marine biodiversity for future generations.”
@@
AltEnergy ”Accelerating the transition to renewable energy to combat climate change and support economic growth through job creation.”, ”Promoting clean and renewable energy to foster economic growth, create jobs, and address climate change effectively.”, ”Public engagement to support and expand carbon-free energy as a crucial step against climate change.”, ”The adoption of wind energy as a crucial, effective solution for reducing CO2 emissions and combating climate change.”, ”Diversified renewable energy solutions to ensure sustainability, grid reliability, and meet increasing demand for clean power across different regions.”, ”Expansion and innovation in bioethanol production and solar energy to support a sustainable, clean energy future.”, ”The positive impact of wind energy, emphasizing job creation, economic growth, environmental benefits, and community engagement.”, ”Political support and action towards clean energy, job creation, climate action, and sustainability initiatives in local communities.”, ”Investment in ICHOR’s revolutionary fuel injection technology to reduce emissions and fuel consumption.”, ”Transitioning from coal to cleaner energy sources to combat climate change and promote sustainability.”, ”The adoption and celebration of renewable energy initiatives while balancing environmental preservation.”
@@
SupportClimatePolicy ”Passing the Build Back Better Act, highlighting its importance for economic growth, job creation, clean energy employment, and combating climate change.”, ”Supporting the Growing Climate Solutions Act and the senators who champion this bipartisan environmental policy.”, ”Supporting the Growing Climate Solutions Act as a means to boost farmer income and combat climate change.”, ”Legislative action and public support for policies that address climate change through clean energy investment and promote social justice.”, ”The passage of the Build Back Better Act, emphasizing its investments in clean energy and climate action.”, ”Halting the Line 3 Pipeline construction to protect the climate, water resources, and Indigenous rights.”, ”Voting for candidates who prioritize climate change action and sustainable policies in their platforms.”, ”The inclusion of bold climate action in legislation, simultaneous with job creation and equity.”
@@
PoliticalAffiliation ”Active political participation and advocacy to influence policies on environmental protection and social justice across various states.”, ”Exposing and combating the corrupt influence of the fossil fuel industry on politics to promote climate action.”, ”Appointments that prioritize environmental protection and resist the fossil fuel industry’s influence in government and finance.”, ”Holding fossil fuel companies accountable through political, public action, and legislative reforms for their environmental impact and misinformation.”, ”Greater scrutiny of political ties to the oil and gas industry and calls for fair regulation.”, ”Understanding and challenging perceived extremism in politics, scrutinizing shifts in political power and ideology, and recognizing ongoing political efforts on issues like infrastructure and climate change.”, ”Supporting candidates whose platforms align with individual voter values and community needs.”, ”Political appointments and decisions that prioritize climate action and reduce the influence of fossil fuels.”
@@
BidenGasPriceIncrease ”Reversing President Biden’s energy policies to support American energy production and reduce gas prices for families.”, ”Against perceived financial mismanagement by ’DC liberals’, linking it to inflation and higher gas prices under Biden’s presidency.”, ”Mobilizing against government mandates perceived to increase gas prices, emphasizing grassroots engagement to protect consumer interests.”, ”Opposing a spending plan criticized for contributing to higher gas prices and economic burdens.”, ”Criticizes President Biden’s economic policies for significantly increasing gas prices and living costs, urging a shift in leadership or policy.”, ”A return to energy independence to control gas prices and criticizes current policies for rising fuel costs.”
@@
AgainstCorpInter ”Supporting Sarah, who prioritizes climate action, honesty, and fiscal responsibility over corporate interests and misinformation.”, ”Grassroots support over corporate interests, emphasizing financial contributions to champion climate, democracy, and economic issues in Iowa’s race.”, ”Grassroots-supported candidates promoting progressive policies over those funded by corporate interests and lobbyists in Congressional races.”, ”Dismantling financial and political support for the fossil fuel industry to favor sustainable, clean energy solutions and accountability.”, ”Electing progressive candidates who prioritize public welfare and grassroots support over corporate interests and profits.”, ”Using the power of voting to challenge corporate interests by supporting policies and candidates focused on public welfare and environmental sustainability.”, ”Empowering grassroots movements over corporate interests, emphasizing small donations to drive significant policy changes for social and environmental justice.”, ”Active resistance against corporate influence in politics and society to achieve environmental, social, and economic reforms.”, ”Grassroots, community-focused leadership challenging established politics and corporate dominance to achieve progressive social and environmental reforms.”, ”Grassroots funding to counteract corporate-funded campaigns, emphasizing the power of collective small donations against big money interests.”
@@
GasTax ”Reassessment of gas tax reliance, emphasizing modernizing revenue models to sustainably fund road maintenance amidst electric vehicle challenges.”, ”Halting proposed gas tax hikes to alleviate financial burdens on residents amidst concerns over rising living costs.”, ”Repealing the CPI portion of the gas tax law to alleviate economic burdens on citizens.”, ”Against implementing a per-mile tax for drivers, citing concerns over increased burdens and privacy issues.”, ”Opposing gas tax increases in Lake and Kane Counties, Illinois, citing high current gas taxes and economic challenges.”, ”Against a proposed fuel mandate in Iowa, framed as a tax on consumers leading to increased gas prices.”, ”Against the implementation of a new gas tax in Connecticut due to its potential financial burdens on families.”, ”The burden of state gas taxes on Californians amid inflation, advocating for measures to ease financial strain on residents.”, ”Lowering gas taxes as part of broader fiscal reforms to reduce the financial burden on citizens.”, ”Emphasizes the need for transparency, accountability, and community involvement in decisions impacting gas station operations and related policies.”
@@
Deforestation ”Roadless protections in the Tongass National Forest to preserve its ecosystem, combat climate change, and support local economies and cultures.”, ”Leveraging consumer purchases for tree planting to combat deforestation and climate change, emphasizing individual contributions to collective environmental efforts.”, ”Global reforestation and individual action against climate change through tree planting and support for sustainable community practices to combat deforestation.”, ”Permanent legislative protection of the Arctic National Wildlife Refuge to preserve biodiversity and prevent environmental degradation from oil and gas extraction.”, ”Preserving urban green spaces and mature trees to maintain air quality, carbon sequestration, and historical landscapes against deforestation for construction projects.”, ”Combating deforestation and climate change through community-driven reforestation efforts and financial support in California.”, ”Stopping the use of forests as fuel for electricity and urging significant climate action to address and mitigate deforestation.”, ”Halting energy projects that cause deforestation, respecting treaties, and protecting ecosystems against corporate environmental damage.”, ”Eliminating conflict palm oil from products to protect rainforests and wildlife, promoting transparency and sustainable sourcing.”, ”Improved forest management and proactive measures like controlled burns to mitigate wildfire risks and prevent deforestation.”, ”Conservation, restoration efforts, and responsible management policies to combat climate change and prevent deforestation.”, ”Restoration and protection of critical habitats to combat deforestation and support ecosystem resilience and biodiversity.”, ”Indigenous-led efforts to protect natural resources and uphold sovereignty as crucial in combating deforestation and environmental degradation.”
@@
Carbon ”The cement and concrete industry’s commitment to environmental sustainability and carbon neutrality to combat climate change by 2050.”, ”Accessible carbon offset solutions, emphasizing affordability, simplicity, and the importance of taking action for Earth Day.”, ”Using Ecologi’s platform to offset carbon footprint, plant trees, and collectively combat climate change.”, ”Implementing carbon pricing strategies through legislative action, communication, and collective efforts to address climate change.”, ”Collaborative efforts across various sectors to reduce carbon emissions and transition towards a sustainable, low-carbon economy.”, ”Innovative technologies and solutions to capture, reduce, and offset carbon emissions for a more sustainable and carbon-neutral future.”, ”Reducing carbon footprints through lifestyle changes, sustainable eating habits, and supporting initiatives aimed at carbon neutrality and environmental sustainability.”

### v3

@@ A.4 Arguments under specific Themes
A.5 Prompting
In this section, we show the prompt templates used in our work to generate a sub-cluster summary (Figure. 5 (a)) and a talking point from each summarized sub-cluster (Figure. 5 (b)). Prompt examples for summarizing top-5 instances and generating talking point from that summary under patriotism theme of climate campaign dataset are shown in Fig. 6 and Fig. 7 respectively.
Figure 6: Prompt example of summarizing top-5 instances under patriotism theme of climate campaign dataset. The black colored segment is the input prompt and the red colored segment is the generated output by the LLMs.
Figure 7: Prompt example of generating talking point from a summary of top-5 instances under patriotism theme of climate campaign dataset. The black colored segment is the input prompt and the red colored segment is the generated output by the LLMs.
A.6 Resulting Arguments
@@ A.6.1 Climate Campaigns
A.6.2 COVID-19 Vaccine Campaigns
A.7 Human Validation
Three NLP and CSS researchers, age range $30$-$45$, $2$ males and $1$ female, did two $1$-hour sessions to validate whether the generated and refined talking points as well as to annotate whether $ads\rightarrow arguments$ mapping is correct or not. The annotators included advanced graduate students and faculty.
A.8 Demographic Analysis
Table 9 shows the most mentioned entities and talking points of targeted ads for young, working-age, older population of CA and TX from climate campaigns dataset.
@@ A.8 Demographic Analysis
Economy_pro "Against natural gas bans due to their potential economic harm to families and workers.", "Joining Energy Citizens to support the natural gas and oil industry’s economic impact and sustainability efforts.", "Acknowledging and supporting the oil and natural gas industry’s critical role in economic growth and job creation.", "Permian Basin’s significant contribution to the energy sector’s growth and the positive economic impact of increased drilling activity.", "Considering the economic and community impact of taxation on the energy sector, emphasizing support for energy-related jobs and services.", "The Line 3 Replacement project’s role in boosting Minnesota’s economy through job creation, supporting Tribal communities, and enhancing safety, while positively impacting the environment.", "Recognizing the oil and gas industry’s vital role in economic growth, while addressing challenges like automation and training the future workforce."
@@
ClimateSolution "Adopting propane as a versatile and environmentally friendly energy solution to support climate change mitigation efforts.", "Leveraging innovative technologies and the synergy between natural gas and renewables as key strategies for climate change mitigation.", "Leveraging diverse, innovative, and market-based strategies to effectively address climate change and promote environmental sustainability.", "Adoption of RGGI in Pennsylvania, international cooperation on climate action, and exploring market-based, conservative solutions for sustainable economic and environmental health.", "Investing in innovative technologies like ICHOR’s RKI Injector and AT&T’s initiatives to drastically reduce emissions and improve fuel efficiency."
@@
Pragmatism "Pragmatic, multi-tiered action on climate change, emphasizing renewable energy investments and policy for sustainable development and job growth.", "Practical investments in sustainable infrastructure and energy diversity to address climate change and ensure future resilience.", "Pragmatic approach to energy transition, emphasizing the need to harmonize environmental sustainability, economic growth, and job preservation.", "Recognizing and utilizing natural gas as a pragmatic, efficient energy solution for reliability and affordability.", "Pragmatic approach to energy policy that balances clean energy transition with economic growth and job preservation.", "Pragmatic approach to energy, balancing economic growth and environmental protection through Colorado’s oil and gas industry.", "Pragmatic approach to energy, emphasizing the need for oil and safe pipeline infrastructure for global energy security.", "Pragmatic policy changes and discussions that address workers’ rights, environmental sustainability, and societal challenges for improvement and growth.", "Pragmatic solutions to environmental challenges through technological innovation, collaboration, and strategic policy efforts to ensure sustainability and security."
@@
Patriotism "Prioritizing domestic energy production as a patriotic duty to ensure America’s independence, security, and economic prosperity.", "Maintaining local oil and gas production to ensure economic and energy security, highlighting patriotism through self-reliance.", "Supporting the oil and gas industry as a patriotic duty, emphasizing its critical role in national security and economic health.", "Supporting natural gas to ensure energy independence and consumer choice, highlighting patriotism through energy security and freedom.", "Rallying around patriotism, energy independence, and conservative values to protect and promote American interests and security.", "Maintaining local oil and gas production to ensure California’s energy independence and protect consumers from rising fuel costs.", "Energy independence for Israel through faith-inspired and scientifically supported oil and gas exploration, and diverse leadership in community engagement."
@@
AgainstClimatePolicy "Renewable energy investment as a strategy for economic growth and job creation while addressing climate change.", "Transitioning to a clean energy economy, emphasizing job creation, sustainability, and combating climate change for economic recovery.", "Significant investment in clean energy infrastructure to create jobs and power homes sustainably, driving economic and environmental progress.", "Passage of the bipartisan infrastructure deal to enhance clean water, transit, and create numerous good jobs.", "American Jobs Plan as a catalyst for economic revitalization through clean energy, creating jobs, and improving infrastructure.", "Enhancing sustainability through clean energy initiatives, education, infrastructure improvement, and worker health and safety in the clean energy sector.", "Local and state-level initiatives that promote clean energy, sustainability, and environmental protection to build a more sustainable future.", "Comprehensive community and environmental betterment through sustainable economic practices, fair wages, and innovative climate solutions.", "Rigorous scrutiny and equitable policies in clean-energy projects to ensure they genuinely benefit the environment and communities."
@@
Economy_clean "Renewable energy investment as a strategy for economic growth and job creation while addressing climate change.", "Transitioning to a clean energy economy, emphasizing job creation, sustainability, and combating climate change for economic recovery.", "Significant investment in clean energy infrastructure to create jobs and power homes sustainably, driving economic and environmental progress.", "Passage of the bipartisan infrastructure deal to enhance clean water, transit, and create numerous good jobs.", "American Jobs Plan as a catalyst for economic revitalization through clean energy, creating jobs, and improving infrastructure.", "Enhancing sustainability through clean energy initiatives, education, infrastructure improvement, and worker health and safety in the clean energy sector.", "Local and state-level initiatives that promote clean energy, sustainability, and environmental protection to build a more sustainable future.", "Comprehensive community and environmental betterment through sustainable economic practices, fair wages, and innovative climate solutions.", "Rigorous scrutiny and equitable policies in clean-energy projects to ensure they genuinely benefit the environment and communities."
@@
HumanHealth "Reducing air pollution from oil and gas industries to protect children’s health and ensure environmental justice.", "Electing knowledgeable leaders and supporting organizations to address climate change’s significant health and humanitarian impacts.", "Climate change’s health impacts through informed initiatives and equity-focused efforts to protect vulnerable communities.", "Leadership that understands and addresses the health impacts of climate change, emphasizing immediate action for global health and stability.", "Environmental health risks, from fracking and gas stoves to substance abuse, to protect and improve public health.", "Combining environmental activism with health equity to combat pollution and support sustainable, just communities.", "Awareness of gas stoves’ health risks, promoting safer alternatives, and implementing preventive measures to protect indoor air quality and prevent asthma.", "Informed, compassionate leadership and global medical support to address healthcare challenges, ensuring dignity and comprehensive care across life stages and crises.", "Addressing and mitigating air pollution sources to protect public health and promote environmental sustainability.", "Proactive engagement and informed responses to health and safety challenges exacerbated by environmental changes and societal issues."
@@
FutureGeneration "Urgent environmental action and policy change to protect future generations from the health impacts of climate change.", "Educating children about climate change through open discussions to inspire future generations to take action for their planet’s health.", "Collective action led by motivated parents to combat climate change for a healthier future for children.", "Empowering the youth to lead climate change solutions, emphasizing education, activism, and support for vulnerable communities.", "Civic engagement and local policy reforms to secure a safer, fairer, and sustainable future for the next generation.", "Sustainable and renewable energy initiatives led by Indigenous and community efforts to ensure a healthier planet for future generations.", "Integrated solutions to climate change that also address social, racial, and economic inequities for future generations’ benefit.", "Sustainable energy and community initiatives that create educational, economic, and environmental benefits for current and future generations."
@@
Environmental "Ending fossil fuel use, implementing clean energy transitions, and enacting protective regulations to combat climate change and pollution.", "Urgent legislative action on climate change to enhance resiliency against extreme weather events and secure a sustainable future.", "Collective action and education to combat climate change, emphasizing community involvement and urgent policy measures.", "Community resilience through local initiatives, transitioning to clean energy, sustainable practices, and comprehensive support for businesses impacted by environmental disasters.", "Environmental protection and Indigenous rights by opposing oil pipelines that threaten natural resources and exacerbate the MMIW epidemic.", "Vigilance and community action against organized crime targeting green energy resources, and promotes sustainable living practices.", "Proactive climate change solutions through innovative forestry, ecosystem restoration projects, clean energy investment, and urgent policy actions to safeguard the future.", "Climate change’s role in exacerbating droughts, fires, and heatwaves, underscored by scientific consensus and observable impacts like Minnesota’s rising temperatures.", "Active participation in environmental conservation efforts led by Derek Kilmer to restore Puget Sound, protect wilderness, and combat climate change."
@@
Animals "Re-protection of gray wolves, recognizing their vital ecological role, and calls for increased conservation efforts and public support.", "Urgent financial contributions, leveraging donation matches to combat threats to endangered wildlife like habitat loss and climate change.", "Urgent climate action to protect Alaska’s fisheries, crucial for both the economy and marine ecosystems.", "Continued investment in climate and clean energy solutions to protect birds from climate change impacts.", "Immediate public action to conserve wildlife and address climate change’s role in accelerating species extinction.", "Habitat restoration and conservation to support wildlife, enhance outdoor activities, and address climate change for future ecological balance.", "Protecting and restoring ocean ecosystems to preserve marine biodiversity and combat climate change’s impacts on animals.", "Awareness and action against the environmental and ethical harms of animal agriculture, emphasizing sustainable choices and animal rights.", "Habitat conservation, creating pollinator gardens, and innovative habitat restoration to protect bats, bees, and wildlife crucial for ecosystem health and biodiversity.", "Promoting renewable energy projects as a means to generate clean power, support local economies, and protect marine biodiversity for future generations."
@@
AltEnergy "Accelerating the transition to renewable energy to combat climate change and support economic growth through job creation.", "Promoting clean and renewable energy to foster economic growth, create jobs, and address climate change effectively.", "Public engagement to support and expand carbon-free energy as a crucial step against climate change.", "The adoption of wind energy as a crucial, effective solution for reducing CO2 emissions and combating climate change.", "Diversified renewable energy solutions to ensure sustainability, grid reliability, and meet increasing demand for clean power across different regions.", "Expansion and innovation in bioethanol production and solar energy to support a sustainable, clean energy future.", "The positive impact of wind energy, emphasizing job creation, economic growth, environmental benefits, and community engagement.", "Political support and action towards clean energy, job creation, climate action, and sustainability initiatives in local communities.", "Investment in ICHOR’s revolutionary fuel injection technology to reduce emissions and fuel consumption.", "Transitioning from coal to cleaner energy sources to combat climate change and promote sustainability.", "The adoption and celebration of renewable energy initiatives while balancing environmental preservation."
@@
SupportClimatePolicy "Passing the Build Back Better Act, highlighting its importance for economic growth, job creation, clean energy employment, and combating climate change.", "Supporting the Growing Climate Solutions Act and the senators who champion this bipartisan environmental policy.", "Supporting the Growing Climate Solutions Act as a means to boost farmer income and combat climate change.", "Legislative action and public support for policies that address climate change through clean energy investment and promote social justice.", "The passage of the Build Back Better Act, emphasizing its investments in clean energy and climate action.", "Halting the Line 3 Pipeline construction to protect the climate, water resources, and Indigenous rights.", "Voting for candidates who prioritize climate change action and sustainable policies in their platforms.", "The inclusion of bold climate action in legislation, simultaneous with job creation and equity."
@@
PoliticalAffiliation "Active political participation and advocacy to influence policies on environmental protection and social justice across various states.", "Exposing and combating the corrupt influence of the fossil fuel industry on politics to promote climate action.", "Appointments that prioritize environmental protection and resist the fossil fuel industry’s influence in government and finance.", "Holding fossil fuel companies accountable through political, public action, and legislative reforms for their environmental impact and misinformation.", "Greater scrutiny of political ties to the oil and gas industry and calls for fair regulation.", "Understanding and challenging perceived extremism in politics, scrutinizing shifts in political power and ideology, and recognizing ongoing political efforts on issues like infrastructure and climate change.", "Supporting candidates whose platforms align with individual voter values and community needs.", "Political appointments and decisions that prioritize climate action and reduce the influence of fossil fuels."
@@
BidenGasPriceIncrease "Reversing President Biden’s energy policies to support American energy production and reduce gas prices for families.", "Against perceived financial mismanagement by ’DC liberals’, linking it to inflation and higher gas prices under Biden’s presidency.", "Mobilizing against government mandates perceived to increase gas prices, emphasizing grassroots engagement to protect consumer interests.", "Opposing a spending plan criticized for contributing to higher gas prices and economic burdens.", "Criticizes President Biden’s economic policies for significantly increasing gas prices and living costs, urging a shift in leadership or policy.", "A return to energy independence to control gas prices and criticizes current policies for rising fuel costs."
@@
AgainstCorpInter "Supporting Sarah, who prioritizes climate action, honesty, and fiscal responsibility over corporate interests and misinformation.", "Grassroots support over corporate interests, emphasizing financial contributions to champion climate, democracy, and economic issues in Iowa’s race.", "Grassroots-supported candidates promoting progressive policies over those funded by corporate interests and lobbyists in Congressional races.", "Dismantling financial and political support for the fossil fuel industry to favor sustainable, clean energy solutions and accountability.", "Electing progressive candidates who prioritize public welfare and grassroots support over corporate interests and profits.", "Using the power of voting to challenge corporate interests by supporting policies and candidates focused on public welfare and environmental sustainability.", "Empowering grassroots movements over corporate interests, emphasizing small donations to drive significant policy changes for social and environmental justice.", "Active resistance against corporate influence in politics and society to achieve environmental, social, and economic reforms.", "Grassroots, community-focused leadership challenging established politics and corporate dominance to achieve progressive social and environmental reforms.", "Grassroots funding to counteract corporate-funded campaigns, emphasizing the power of collective small donations against big money interests."
@@
GasTax "Reassessment of gas tax reliance, emphasizing modernizing revenue models to sustainably fund road maintenance amidst electric vehicle challenges.", "Halting proposed gas tax hikes to alleviate financial burdens on residents amidst concerns over rising living costs.", "Repealing the CPI portion of the gas tax law to alleviate economic burdens on citizens.", "Against implementing a per-mile tax for drivers, citing concerns over increased burdens and privacy issues.", "Opposing gas tax increases in Lake and Kane Counties, Illinois, citing high current gas taxes and economic challenges.", "Against a proposed fuel mandate in Iowa, framed as a tax on consumers leading to increased gas prices.", "Against the implementation of a new gas tax in Connecticut due to its potential financial burdens on families.", "The burden of state gas taxes on Californians amid inflation, advocating for measures to ease financial strain on residents.", "Lowering gas taxes as part of broader fiscal reforms to reduce the financial burden on citizens.", "Emphasizes the need for transparency, accountability, and community involvement in decisions impacting gas station operations and related policies."
@@
Deforestation "Roadless protections in the Tongass National Forest to preserve its ecosystem, combat climate change, and support local economies and cultures.", "Leveraging consumer purchases for tree planting to combat deforestation and climate change, emphasizing individual contributions to collective environmental efforts.", "Global reforestation and individual action against climate change through tree planting and support for sustainable community practices to combat deforestation.", "Permanent legislative protection of the Arctic National Wildlife Refuge to preserve biodiversity and prevent environmental degradation from oil and gas extraction.", "Preserving urban green spaces and mature trees to maintain air quality, carbon sequestration, and historical landscapes against deforestation for construction projects.", "Combating deforestation and climate change through community-driven reforestation efforts and financial support in California.", "Stopping the use of forests as fuel for electricity and urging significant climate action to address and mitigate deforestation.", "Halting energy projects that cause deforestation, respecting treaties, and protecting ecosystems against corporate environmental damage.", "Eliminating conflict palm oil from products to protect rainforests and wildlife, promoting transparency and sustainable sourcing.", "Improved forest management and proactive measures like controlled burns to mitigate wildfire risks and prevent deforestation.", "Conservation, restoration efforts, and responsible management policies to combat climate change and prevent deforestation.", "Restoration and protection of critical habitats to combat deforestation and support ecosystem resilience and biodiversity.", "Indigenous-led efforts to protect natural resources and uphold sovereignty as crucial in combating deforestation and environmental degradation."
@@
Carbon "The cement and concrete industry’s commitment to environmental sustainability and carbon neutrality to combat climate change by 2050.", "Accessible carbon offset solutions, emphasizing affordability, simplicity, and the importance of taking action for Earth Day.", "Using Ecologi’s platform to offset carbon footprint, plant trees, and collectively combat climate change.", "Implementing carbon pricing strategies through legislative action, communication, and collective efforts to address climate change.", "Collaborative efforts across various sectors to reduce carbon emissions and transition towards a sustainable, low-carbon economy.", "Innovative technologies and solutions to capture, reduce, and offset carbon emissions for a more sustainable and carbon-neutral future.", "Reducing carbon footprints through lifestyle changes, sustainable eating habits, and supporting initiatives aimed at carbon neutrality and environmental sustainability."